a right circular cylinder has a volume of 1000 cubic inches and a height of 8 inches. what is the radius?
Answer by mananth(16946)

The volume of a cylinder is given by pi*r^2 *h
1000= 3.14*r^2*8
r^2= 1000/3.14*8
r^2=39.8
r=6.3 inches